A person with a mass of 90 kg contains 60 kg of water. How many kilograms of water are in a person with a mass of 50 kg?
Answer by Boreal(15235)

90 is to 60 as 50 is to x
90/60=50/x
cross-multiply
90x=3000
x=33 (1/3) kg.